A COMPLETE SPECTROSCOPIC MAP AND NARROW-BAND IMAGING OF SMALL PAHS IN THE RED RECTANGLE NEBULA

\underline{UMA P. VIJH}, ADOLF N. WITT, Ritter Astrophysical Research Center, Univeristy of Toledo, Toledo, OH 43606; DONALD G. YORK, VIKRAM DWARKADAS, University of Chicago, Chicago, IL 60637; BRUCE E. WOODGATE, NASA Goddard Space Flight Center, Greenbelt, MD 20771; POVILAS PALUNAS, McDonald Observatory, University of Texas at Austin, Austin, TX 78712.

Electronic flurescence spectra (peak ∼ 375 nm) in reflection nebulae have helped to identify the largest molecules that have been detected so far in the interstellar medium: 3-4 ringed polycyclic aromatic hydrocarbons (PAHs). This detection of blue luminescence (BL) by small, neutral PAHs was first made in the peculiar, proto-planetary nebula, the Red Rectangle. This first detection and subsequent observations in other reflection nebulae reveal spatial variations in the BL spectrum indicating a change in the size distribution/ionisation state of the emitters. Data from an ongoing, complete spectroscopic survey and narrow-band imaging of the Red Rectangle will be presented. This study sheds light on the spatial distribution, ionization state and the size distributions of the small PAHs in this nebula.